# Roadside Units Assisted Localized Automated Vehicle Maneuvering: An Offline Reinforcement Learning Approach

### v1

@@ I Introduction
Traffic intersections are one of the most hazardous areas in transportation systems due to the frequent presence of blind spots, vulnerable pedestrians, and stop-and-go traffic [1]. A report from the US Federal Highway Administration (FHA) reveals that intersections are hotspots for vehicular accidents, with 40% of all traffic crashes and an alarming 50% of serious collisions occurring within these areas [2]. Despite the advent and increasing adoption of connected and automated vehicles (CAVs), which promise to enhance traffic safety by minimizing human error, ensuring the safety of CAVs at intersections, particularly those without signalization, remains a formidable challenge [3].
@@
With recent advances in vehicle-to-everything (V2X) technologies, autonomous intersection control (AIC) has emerged as a promising solution for enhancing road safety and improving traffic efficiency [4]. Today’s V2X technology encompasses a suite of communication links: vehicle-to-vehicle (V2V), vehicle-to-infrastructure (V2I), vehicle-to-pedestrian (V2P), and vehicle-to-cloud (V2C) [5]. Notably, V2I communication facilitates real-time data exchange between CAVs and smart roadside infrastructure, laying the groundwork for intelligent transportation ecosystems [6]. This synergy between CAVs and roadside infrastructure aids in mitigating the inherent risks associated with traffic intersections by leveraging shared data to optimize driving decisions.
In light of these advancements, there is an increasing focus on integrating roadside infrastructure functionalities, e.g., smart traffic signals and roadside units (RSUs), into the CAV operational frameworks. This integration can be broadly categorized into two key areas: cooperative perception and cooperative driving. Cooperative perception extends the sensory capabilities of vehicles, enabling them to detect potential hazards and traffic conditions beyond their sensing views [7]. For example, in [8], a framework of blind-spot visualization to see through obstacles was proposed for safe driving via millimeter-wave communications. In [9], a multi-camera data fusion system was employed to realize vision-based cooperative perception. Meanwhile, cooperative driving has also emerged as a critical area of interest within smart mobility and intelligent transportation systems (ITS). Particularly, authors in [10, 11, 12] highlighted the importance of driving strategies in the context of traffic intersections, aiming to reduce energy consumption, enhance traffic safety, and improve traveling efficiency, respectively. It is worth noting that the distinct characteristics of traffic intersections, such as their road topology, the volume of traffic, and the diversity of traffic participants have not been considered in the existing literature. These features mandate the development of localized driving strategies tailored to each specific intersection scenario. Such a research gap points to a pressing need for a more sophisticated and intersection-specific methodology for formulating CAV operational strategies.
@@
In contrast to these prior works, the main contribution of this paper is a novel RSU-assisted cooperative maneuvering system that leverages offline reinforcement learning (RL) to enhance the operational strategies of CAVs at intersections. Our approach uniquely integrates the edge computing capabilities of RSUs [13] with an innovative offline RL framework to create a system that not only learns from historical traffic data but also adapts to the diverse characteristics of different intersections. This methodology enables the development of highly localized and effective driving strategies that account for unique configurations and dynamic traffic environments of each intersection. Simulation results demonstrate that the proposed system offers a substantial enhancement in traffic safety and efficiency within RSU-installed areas, outperforming conventional autonomous driving systems. In summary, our key contributions include:
@@
  • System design: Conceptual design of a novel system that utilizes RSUs to assist in the maneuvering of CAVs at intersections is proposed, by leveraging edge computing and V2X communications to enable object detection and cooperative perception.
@@
  • Intersection-specific strategies: Based on the system design, we facilitate localized driving strategies that are tailored to the unique configurations and dynamic traffic conditions of individual intersections.
@@
  • Offline RL: We deploy an innovative offline RL framework that capitalizes on historical traffic data to optimize driving strategies for CAVs.
@@
  • Safety and efficiency improvements: The proposed system demonstrates significant improvements in traffic efficiency in areas with RSU installations while ensuring safety measures, as evidenced by comparative simulations with conventional autonomous driving systems.

### v2

@@ I Introduction
Traffic intersections are one of the most hazardous areas in transportation systems due to frequent blind spots, vulnerable pedestrians, and stop-and-go traffic [1]. A report from the US Federal Highway Administration (FHA) reveals that intersections are hotspots for vehicular accidents, with 40% of all traffic crashes and 50% of serious collisions occurring in these areas [2]. Despite the advent and increasing adoption of connected and automated vehicles (CAVs), ensuring their safety at intersections, particularly those without signalization, remains a formidable challenge [3].
@@
Recent advances in vehicle-to-everything (V2X) technologies have given rise to vehicle-road collaboration, which holds promise for enhancing road safety and traffic efficiency [4]. V2X technology includes vehicle-to-vehicle (V2V), vehicle-to-infrastructure (V2I), vehicle-to-pedestrian (V2P), and vehicle-to-cloud (V2C) communication [5]. Notably, V2I communication enables real-time data exchange between CAVs and smart roadside infrastructure, laying the groundwork for intelligent transportation ecosystems [6]. This synergy between CAVs and roadside infrastructure aids in mitigating the inherent risks associated with traffic intersections by leveraging shared data to optimize driving decisions.
In light of these advancements, there is an increasing focus on integrating roadside infrastructure functionalities, e.g., smart traffic signals and roadside units (RSUs), into the CAV operational frameworks. This integration can be broadly categorized into two key areas: cooperative perception and cooperative driving. Cooperative perception extends the sensory capabilities of vehicles, enabling them to detect potential hazards and traffic conditions beyond their sensing views [7]. On the other hand, cooperative driving allows CAVs to coordinate their maneuvers in complex environments like traffic intersections and highways, thereby improving overall traffic safety and efficiency.
Intersection-tailored cooperative perception and driving have witnessed significant advancements in recent years. On the one hand, various studies have focused on enhancing CAV operations by using cooperative perception [8, 9, 10, 11, 12, 13, 14, 15, 16]. For example, authors in [11] propose a birds-eye-view grid occupancy prediction method to support various applications like driving safety warning and cooperative traffic control. In [12], a perception task-oriented information-sharing network is developed to enable occlusion-free environmental awareness for CAVs. Especially, the work in [13, 14] also notices the importance of unique features of different intersections, where intersection-specific strategies are deployed for driver-intended movements prediction and road-to-vehicle visual information sharing, respectively. On the other hand, research in cooperative driving mainly focuses on developing sophisticated algorithms and robust communication protocols that enhance the collaborative decision-making capabilities of CAVs in dynamic and uncertain environments [17, 18, 19, 20]. For instance, multi-vehicle coordination systems utilize real-time data sharing and predictive modeling to enable vehicles to anticipate the actions of others and adjust their trajectories accordingly, thereby reducing the risk of collisions and improving traffic flow [17, 18, 19]. Advanced intersection management models integrate traffic signal optimization and vehicle trajectory planning, considering factors such as speed, density, and pedestrian movements, to minimize congestion and enhance safety [20].
Most existing research in cooperative perception and cooperative driving focuses on sharing dynamic information like vehicle states and predicted behaviors. However, the integration of static information and inherent road traffic features to develop local strategies can also significantly advance autonomous driving. For example, a human driver commuting daily past an intersection near a kindergarten will quickly learn to pay extra attention and adopt a more careful and conservative driving strategy. How to enable CAVs to gain such knowledge is the core motivation of this research.
@@
In this paper, we present a novel RSU-assisted cooperative maneuvering system that leverages offline reinforcement learning (RL) to enhance CAV operational strategies at intersections. Our approach uniquely integrates the edge computing capabilities of RSUs [21] with an innovative offline RL framework to create a system that not only learns from historical traffic data but also adapts to the diverse characteristics of different intersections. This methodology enables the development of highly localized and effective driving strategies that account for unique configurations and dynamic traffic environments of each intersection. Simulation results demonstrate that the proposed system offers a substantial enhancement in traffic safety and efficiency within RSU-installed areas, outperforming conventional autonomous driving systems. In summary, our key contributions include:
@@
  • System Design: We propose a conceptual design of a novel system that uses RSUs to assist CAV maneuvering at intersections by leveraging edge computing and V2X communications for object detection and cooperative perception.
@@
  • Intersection-Specific Strategies: We facilitate localized driving strategies tailored to the unique configurations and dynamic traffic conditions of individual intersections based on the system design.
@@
  • Offline RL: We deploy an innovative offline RL framework that utilizes historical traffic data to optimize driving strategies for CAVs.
@@
  • Safety and Efficiency Improvements: The proposed system demonstrates significant improvements in traffic efficiency and safety in RSU-installed areas, as evidenced by comparative simulations with conventional autonomous driving systems.